# SALM: A Multi-Agent Framework for Language Model-Driven Social Network Simulation

### v1

@@ 1.1 Research Contributions
Our work makes several theoretical and methodological contributions to the field of computational social science. We develop a novel synthesis between LLMs and traditional social simulation methods, grounded in theories of communicative action Habermas (1984) and network-based conceptualizations of social network formation relevant to observable interaction patterns Wasserman and Faust (1994). While broader sociological theories of tie formation in social networks exist Bourdieu (1986), our focus is on its operationalization within network dynamics. By leveraging recent advances in zero-shot reasoning Kojima et al. (2022) and adaptive language modeling Hu et al. (2021), we demonstrate how careful integration of LLMs can enable more nuanced modeling of social phenomena while maintaining computational feasibility. Our approach builds on emerging understandings of LLM capabilities Bommasani et al. (2022) while addressing key challenges in long-term behavioral consistency De Lange et al. (2021). This provides ABM practitioners with a framework capable of simulating more complex and adaptive agent behaviors over longer durations than previously feasible.
@@ 2.1 Agent-Based Social Modeling
Traditional approaches to social simulation have relied on rule-based systems Gilbert and Troitzsch (2005) and statistical models Barabási and Albert (1999). While valuable, rule-based systems often require extensive specification and struggle to adapt to unforeseen contexts, limiting the representation of complex decision-making and ’natural’ language-based interaction. Recent work has demonstrated remarkable advances in generating more believable and contextually adaptive human behavior using language models Park et al. (2023), moving beyond rigid rules. Frameworks like OASIS Mehta et al. (2023) show how online adaptation enables more natural social interactions, while S3 Gao et al. (2023) introduces sophisticated agent architectures specifically designed for social network simulation. Recent innovations in multiagent frameworks Yang et al. (2023) and meta-programming approaches Zhang et al. (2023) have further expanded the capabilities of LLM-based social simulation. However, these systems face significant challenges in computational scalability and long-term coherence Park et al. (2024), which SALM addresses through its hierarchical prompting and memory management.
@@ 3.3 Interaction Dynamics
Agent interactions follow a sophisticated protocol that balances emotional influence with behavioral consistency:
where represents emotional state (e.g., using PAD vectors encoding Pleasure, Arousal, Dominance), is the interaction impact (e.g., scaled by the emotional valence and social importance of the interaction), is the cognitive context (e.g., relevance of the interaction to agent goals), is a natural decay factor, is the emotional influence coefficient, quantifying the impact of recent interactions on the agent’s emotional state, while is the cognitive context coefficient, representing the weight of cognitive or goal-related factors in emotional updating. These coefficients are set based on empirical tuning to balance the influence of social and cognitive factors in agent emotion dynamics. This mechanism ensures stable emotional dynamics while preserving individual personality traits, with drift bounded by:
@@ 3.4 Memory Management
where represents short-term active memory, the cache layer and long-term storage, implementing full attention processing for immediate context, sparse attention with importance sampling for recent history, and compressed storage with similarity-based retrieval for long-term memory respectively. The system computes retrieval scores using:
where represents the importance factor computed through gradient-based memory consolidation. This architecture achieves remarkable efficiency in multi-agent scenarios, with cache hit rates reaching 80% while maintaining high context relevance. Moreover, the system demonstrates exceptional scalability, with memory growth constrained to just 9.5% between timesteps 1000-4000 during extended simulations.
@@ 4 Validation and Evaluation
The Stanford SNAP ego networks dataset used in our experiments is publicly available at https://snap.stanford.edu/data/ego-Facebook.html. We use the Facebook, Google+, and Twitter ego networks for validation.
Although the SNAP ego networks are static snapshots, we validate long-term social interaction modeling by running extended simulations on these fixed network structures. Agents interact, update memory, and evolve over thousands of timesteps, allowing us to assess stability, behavioral drift, and emergent phenomena over time. While this approach does not capture real-world network evolution, it provides a controlled environment for evaluating persistent agent behavior and long-term coherence. Limitations of this methodology are discussed in Section 4.5.
Metric | T | F | G+ | Notes |
Echo Chamber | .89 | .84 | .91 | Observed polarization |
Info. Cascades | .82 | .79 | .88 | Viral spread match |
Group Polarization | .87 | .91 | .85 | Opinion shift agree |
Comm. Resilience | .85 | .88 | .83 | Group survival rate |
Interaction Dyn. | .81 | .86 | .89 | Temporal patterns |
Network Growth | .79 | .83 | .87 | Growth correlation |
@@ 4.1 Real-World Validation Methodology
At the micro level, the MicroValidator analyzed individual agent behaviors and social tie formation patterns, assessing metrics like interaction frequency distributions, reciprocity rates, and the propensity for triadic closure following interactions. Interaction quality was quantified using , where measures the quality between agents and at time . This achieved correlation coefficients of r = 0.86 ± 0.04 for interaction dynamics against empirical benchmarks derived from the datasets. This granular analysis employed pattern matching algorithms to evaluate behavioral consistency and interaction quality metrics.
@@
The meso-level MesoValidator examined community structures and group dynamics, with particular emphasis on replicating empirically observed patterns of social circle formation (e.g., community size distributions, internal density) and their structural properties. Group dynamics were assessed via , where is the set of communities, structural cohesion, and emotional coherence. Community detection employed the Louvain method with resolution parameter optimization, achieving correlations of 0.84-0.91 (Table 1) for structural pattern replication against the ground-truth communities in the Facebook dataset. The framework demonstrated exceptional fidelity in replicating empirical community patterns through hierarchical comparison metrics.
@@
At the macro level, our MacroValidator achieved precise replication of global network properties evident in the datasets, including clustering coefficient (model: 0.31 vs Facebook: 0.31), average path length (model: 4.7 vs Facebook: 4.7), and degree distribution power-law exponent (, matching typical social network findings). Network-level properties were evaluated using , where represents structural metrics, degree distribution, and evolutionary dynamics. The coefficients , , and are weighting factors that determine the relative importance of structure, distribution, and dynamics, respectively, in the overall macro validation score; these were tuned empirically. Network evolution tracking demonstrated stable structural metrics throughout extended simulations (see Section 4.2). While structural replication was the primary focus here, the feature-level validation below assesses behavioral consistency at the agent level.
@@
Feature-level validation employed personality vector comparison between simulated agents and derived features from real user data (where available and anonymized in datasets), achieving mean correlation of 0.88 () across platforms. The validation protocol employed rigorous statistical analysis at each tier, with all reported correlations achieving statistical significance () across multiple independent simulation runs (). Validation simulations were typically initialized with agent characteristics sampled from empirical distributions or set to neutral defaults, then run for up to 4000 timesteps. At each step, agents perceived their environment, updated memory and emotional state, generated actions via the LLM, and executed valid actions, leading to network and state evolution. This comprehensive approach enables systematic evaluation of both structural and behavioral fidelity while maintaining computational efficiency through our hierarchical caching architecture.
@@ 4.4 Ablation Studies
Network-Relationship Deterioration: When memory mechanisms are compromised, agents lose track of their online social history. This leads to observable errors like agents sending repeated friend requests to existing connections, ”forgetting” previous online interactions (e.g., asking the same introductory questions in DMs), or failing to recognize members of shared online communities. Consequently, the simulated social network fails to develop realistic structures; online tie strengths decay arbitrarily, online community cohesion weakens, and the network may fragment unrealistically as agents cannot maintain or build upon past digital interactions.
@@ A.1 Personality Drift Bounds
For any agent with personality vector at time , the drift after interactions is bounded by:
where is the learning rate and is the context sensitivity factor.
@@
With our adaptive learning rate schedule:
Summing the series, we obtain:
Using the approximation for large t:
Thus, the bound becomes:
For large k, this simplifies to:
@@
For each timestep t: 1) New prompts are generated with probability 2) Cache hits occur with probability
@@
This proof provides theoretical justification for our observed 73% reduction in token usage and 80% cache hit rates, as the logarithmic growth in cache size enables efficient reuse of common interaction patterns while maintaining behavioral diversity.

### v2

@@ 1.1 Research Contributions
Our work makes several theoretical and methodological contributions to the field of computational social science. We develop a novel synthesis between LLMs and traditional social simulation methods, grounded in theories of communicative action Habermas (1984) and network-based conceptualizations of social network formation relevant to observable interaction patterns Wasserman and Faust (1994). While broader sociological theories of tie formation in social networks exist, our focus is on its operationalization within network dynamics. By leveraging recent advances in zero-shot reasoning Kojima et al. (2022) and adaptive language modeling Hu et al. (2021), we demonstrate how careful integration of LLMs can enable more nuanced modeling of social phenomena while maintaining computational feasibility. Our approach builds on emerging understandings of LLM capabilities Bommasani et al. (2022) while addressing key challenges in long-term behavioral consistency De Lange et al. (2021). This provides ABM practitioners with a framework capable of simulating more complex and adaptive agent behaviors over longer durations than previously feasible.
@@ 2.1 Agent-Based Social Modeling
Traditional approaches to social simulation have relied on rule-based systems Gilbert and Troitzsch (2005) and statistical models Barabási and Albert (1999). While valuable, rule-based systems often require extensive specification and struggle to adapt to unforeseen contexts, limiting the representation of complex decision-making and “natural” language-based interaction. Recent work has demonstrated remarkable advances in generating more believable and contextually adaptive human behavior using language models Park et al. (2023), moving beyond rigid rules. Frameworks like OASIS Mehta et al. (2023) show how online adaptation enables more natural social interactions, while S3 Gao et al. (2023) introduces sophisticated agent architectures specifically designed for social network simulation. Recent innovations in multiagent frameworks Yang et al. (2023) and meta-programming approaches Zhang et al. (2023) have further expanded the capabilities of LLM-based social simulation. However, these systems face significant challenges in computational scalability and long-term coherence Park et al. (2024), which SALM addresses through its hierarchical prompting and memory management.
@@ 3 Model Architecture and Operation
3.1 Notation
| Symbol | Meaning |
| Personality vector of agent | |
| Emotional state at time (PAD) | |
| Network position features at time | |
| / | Retrieval / update context (e.g., ) |
| Similarity function (cosine, range ) | |
| Deterministic normalization mapping | |
| Hash function for cache keys | |
| Goal, action, relationship factors |
@@ 3.3 Language Model Integration Architecture
| Step | Description | Variables |
| 1 | Context Gathering | |
| 2 | Emotional Update | |
| 3 | Decision Generation | |
| 4 | Action Execution | network update |
| 5 | Memory Update |
@@ 3.4 Interaction Dynamics
Agent interactions follow a sophisticated protocol that balances emotional influence with behavioral consistency, with
where represents emotional state (e.g., using PAD vectors encoding Pleasure, Arousal, Dominance), is the interaction impact (e.g., scaled by the emotional valence and social importance of the interaction), is the cognitive context (e.g., relevance of the interaction to agent goals), is a natural decay factor, is the emotional influence coefficient quantifying the impact of recent interactions on the agent’s emotional state, and is the cognitive context coefficient representing the weight of cognitive or goal-related factors in emotional updating. These coefficients are set based on empirical tuning to balance the influence of social and cognitive factors in agent emotion dynamics. This mechanism ensures stable emotional dynamics while preserving individual personality traits, with drift bounded by:
@@ 3.5 Memory Management
where represents short-term active memory, the cache layer and long-term storage, implementing full attention processing for immediate context, sparse attention with importance sampling for recent history, and compressed storage with similarity-based retrieval for long-term memory respectively. The system computes retrieval scores using
where is the retrieval context (e.g., ), denotes cosine similarity, and is an importance weight. This architecture achieves remarkable efficiency in multi-agent scenarios, with cache hit rates reaching 80% while maintaining high context relevance. Moreover, the system demonstrates exceptional scalability, with memory growth constrained to just 9.5% between timesteps 1000–4000 during extended simulations.
@@ 4 Validation and Evaluation
The Stanford SNAP ego networks dataset used in our experiments is publicly available (Facebook: https://snap.stanford.edu/data/ego-Facebook.html, Google+: https://snap.stanford.edu/data/ego-Gplus.html, Twitter: https://snap.stanford.edu/data/egonets-Twitter.html). We use the Facebook, Google+, and Twitter ego networks for validation. Figure 2 compares key network metrics between our model and Facebook ego networks over time.
Although the SNAP ego networks are static snapshots, we validate long-term social interaction modeling by running extended simulations on these fixed network structures. Agents interact, update memory, and evolve over thousands of timesteps, allowing us to assess stability, behavioral drift, and emergent phenomena over time. While this approach does not capture real-world network evolution, it provides a controlled environment for evaluating persistent agent behavior and long-term coherence. Limitations of this methodology are discussed in Section 4.6, and external validity is considered in the paragraph below.
External Validity.
Our stability results should be interpreted as behavioral stability under extended interaction horizons on fixed topology. In dynamic real-world networks, exogenous tie formation and dissolution can alter opportunities and constraints. We therefore view the static-graph evaluation as isolating agent-level stability from topological drift. Future work will evaluate SALM on dynamic graphs (e.g., time-sliced ego networks) to measure robustness when both preferences and topology co-evolve.
Metric T F G+ Notes Echo Chamber .89 .84 .91 Observed polarization Info. Cascades .82 .79 .88 Viral spread match Group Polarization .87 .91 .85 Opinion shift agree Comm. Resilience .85 .88 .83 Group survival rate Interaction Dyn. .81 .86 .89 Temporal patterns Network Growth .79 .83 .87 Growth correlation
@@ 4.1 Real-World Validation Methodology
At the micro level, the MicroValidator analyzed individual agent behaviors and social tie formation patterns, assessing metrics like interaction frequency distributions, reciprocity rates, and the propensity for triadic closure following interactions. Interaction quality was quantified using , where measures the quality between agents and at time . Here, is computed as cosine similarity between action embeddings modulated by valence alignment and normalized reciprocity, providing a bounded proxy for dyadic interaction quality. This achieved correlation coefficients of r = 0.86 ± 0.04 for interaction dynamics against empirical benchmarks derived from the datasets. This granular analysis employed pattern matching algorithms to evaluate behavioral consistency and interaction quality metrics.
@@
The meso-level MesoValidator examined community structures and group dynamics, with particular emphasis on replicating empirically observed patterns of social circle formation (e.g., community size distributions, internal density) and their structural properties. Group dynamics were assessed via , where is the set of communities, denotes structural cohesion (e.g., normalized internal density in ), and denotes emotional coherence (e.g., inverse variance of group PAD trajectories rescaled to ). Community detection employed the Louvain method blondel:louvain2008 with resolution parameter optimization, achieving correlations of 0.84–0.91 (Table 3) for structural pattern replication against the ground-truth communities in the Facebook dataset. The framework demonstrated exceptional fidelity in replicating empirical community patterns through hierarchical comparison metrics.
@@
At the macro level, our MacroValidator achieved precise replication of global network properties evident in the datasets, including clustering coefficient (model: 0.31 vs Facebook: 0.31), average path length (model: 4.7 vs Facebook: 4.7), and degree distribution power-law exponent (, matching typical social network findings). Network-level properties were evaluated using , where represents structural metrics, degree distribution, and evolutionary dynamics. Coefficients satisfy and (empirically tuned). Network evolution tracking demonstrated stable structural metrics throughout extended simulations (see Section 4.2). While structural replication was the primary focus here, the feature-level validation below assesses behavioral consistency at the agent level.
4.2 Reproducibility Checklist
We aim to maximize reproducibility and ease of adoption. Our artifact includes: (1) complete code and configuration files (repository link provided in the arXiv version), (2) exact prompts and cache configurations, (3) dataset access instructions for SNAP ego networks, (4) scripts for deterministic runs with fixed seeds and number of trials, and (5) hardware and runtime details for each experiment (including steps, tokens, RAM, and wall-clock time).
@@ 4.2 Reproducibility Checklist
Feature-level validation employed personality vector comparison between simulated agents and derived features from real user data (where available and anonymized in datasets), achieving mean correlation of 0.88 () across platforms. The validation protocol employed rigorous statistical analysis at each tier, with all reported correlations achieving statistical significance () across multiple independent simulation runs ().
Behavioral Coherence Metric.
We define coherence as the Spearman correlation between predicted agent-level behavior features and ground-truth proxies, averaged across features and agents, rescaled to . Concretely, for feature set , agent set , and runs , , where is Spearman’s computed per run and aggregates over agents.
Statistical Reporting.
Unless otherwise noted, we report: sample size , 95% confidence intervals from nonparametric bootstrap over runs (), test type (paired t-test or Wilcoxon), number of independent runs , and fixed random seeds. Seeds, splits, and configs are released for full reproducibility. Validation simulations were typically initialized with agent characteristics sampled from empirical distributions or set to neutral defaults, then run for up to 4000 timesteps. At each step, agents perceived their environment, updated memory and emotional state, generated actions via the LLM, and executed valid actions, leading to network and state evolution. This comprehensive approach enables systematic evaluation of both structural and behavioral fidelity while maintaining computational efficiency through our hierarchical caching architecture.
@@ 4.5 Ablation Studies
Ablation Protocol.
All ablations use identical datasets, agent initializations, seeds, and evaluation windows as the full model. We vary one component at a time and report mean 95% CI over independent runs; full configs and seeds are released.
@@
Network-Relationship Deterioration: When memory mechanisms are compromised, agents lose track of their online social history. This leads to observable errors like agents sending repeated friend requests to existing connections, “forgetting” previous online interactions (e.g., asking the same introductory questions in DMs), or failing to recognize members of shared online communities. Consequently, the simulated social network fails to develop realistic structures; online tie strengths decay arbitrarily, online community cohesion weakens, and the network may fragment unrealistically as agents cannot maintain or build upon past digital interactions.
@@ A.1 Personality Drift Bounds
For any agent with personality vector at time , under bounded gradients and a diminishing learning rate schedule with , the drift after interactions is bounded by
@@
With our diminishing learning rate schedule and bounded gradients, we have
Absorbing constants into yields the stated logarithmic bound.
@@
For each timestep , we assume:
1.
New prompts are generated with probability .
2.
Cache hits occur with probability .
@@
This proof provides theoretical justification for our observed 73% reduction in token usage and 80% cache hit rates, as the logarithmic growth in cache size enables efficient reuse of common interaction patterns while maintaining behavioral diversity. For agent identity over long horizons, our bounded drift result aligns conceptually with recent persona consistency evaluations in LLMs serrano:persona2024; SALM’s logarithmic drift bound offers a formal counterpart to empirical persona stability findings.